A chemist needs 110 milliters of a 22% solutions but only has 10% and 54% solutions available. Find how many milliters of each solution that should be mixed to get the desired solution?
I don't know how to set this up

Answer by checkley77(12844)

.54X+.10(110-X)=.22*110
.54X+11-.10X=24.2
.44X=24.2-11
.44X=13.2
X=13.2/.44
X=30 ML. OF 54% SOLUTION IS USED.
110-30=80 ML. OF 10% SOLUTION IS USED.
PROOF:
.54*30+.10*80=.22*110
16.2+8=24.2
24.2=24.2